Given the supply function and inverse demand function, Qs = 40 + 20 P and p=10-1/10Q

respectively. Determine the equilibrium price and quantity?


1
Expert's answer
2021-04-29T07:35:28-0400

"\frac{40-Q}{20}=10-\frac{1}{10Q}"

"\frac{40-Q}{20}=10-0.1Q"

"\frac{40-Q-20(10-0.1Q)}{20}=0"

"\frac{-160+2Q}{20}=0"

Q=80

"Qs = 40 + 20 P"

"80= 40 + 20 P"

P=2

the equilibrium price and quantity is 2 and 80